A company has common stock that can be sold for $33.09 per share. The stock paid a dividend at the end of last year of $1.47. Dividends are expected to grow at an annual rate of 8% indefinitely. Flotation costs associated with the sale of stock equal $5.68 per share. What is the corporation’s cost of external equity? Submit your answer as a percentage and round to two decimal places (Ex. 0.00%)

Carnet Corp. will finance its next major expansion with 40% debt, 10% preferred stock, and 50% new common stock. Carnet’s after-tax cost of debt is 3.8%, cost of preferred stock is 6.8%, cost of retained earnings is 13.4%, and cost of new common stock is 19.9%. What is the corporation’s weighted average cost of capital? Submit your answer as a percentage and round to two decimal places (Ex. 0.00%)
Answer #1

1)

Price = 33.09 - 5.68 = 27.41

Year 1 dividend = 1.47 (1 + 8%) = 1.5876

Cost of external equity = (D1 / price) + growth rate

Cost of external equity = (1.5876 / 27.41) + 0.08

Cost of external equity = 0.05792 + 0.08

Cost of external equity = 0.1379 or 13.79%

2)

Weighted average cost of capital = Weights * costs

Weighted average cost of capital = 0.4*0.038 + 0.1*0.068 + 0.5*0.199

Weighted average cost of capital = 0.0152 + 0.0068 + 0.0995

Weighted average cost of capital = 0.1215 or 12.15%